EXPERIMENTAL STUDIES OF THE ASYMMETRICALLY DEUTERATED METHOXY RADICALS WITH THE FASSST TECHNIQUE

\underline{DMITRY G. MELNIK}, VADIM STAKHURSKY, VLADIMIR A. LOZOVSKY, TERRY A. MILLER and C. BRADLEY MOORE, The Ohio State University, Dept. of Chemistry, Laser Spectroscopy Facility, 120 W. 18th Avenue, Columbus, Ohio 43210; FRANK C. DE LUCIA, The Ohio State University, Dept. of Physics, Microwave Laboratory, 174 W. 18th Avenue, Columbus OH 43210.

Asymmetrically deuterated methoxy radicals were experimentally studied in a direct absorption experiment using the pulsed jet FASSST spectrometer^{a} in the range of 153-251 GHz . The radical species were produced in a gas discharge in a supersonic jet by a Stark-free coaxial electrode arrangement, mounted on a top of the nozzle. Observed transitions were attributed to methoxy only if they satisfied the following criteria:

i) originated from a transient species, observable only under discharge conditions; ii) appeared in spectra recorded with both precursors used (methanol and methyl nitrite); iii) matched no reported transitions in other molecules (e.g. formaldehyde) found in discharge; iv) exhibited paramagnetic properties. The spectra were assigned by comparison of the experimentally obtained pattern with ones produced by the SpecView program using the available^{b} Hamiltonian and molecular constants with a subsequent fit of the experimental data to the theoretical model.

aD.Melnik, S.Gopalakrishnan, T.A.Miller, S.P.Belov and F.C.De Lucia, J.Chem.Phys., 114, 6100 (2001)

bI.Kalinovki, “Laser Induced Fluorescence Spectroscopy of CHD_{2}O and CH_{2}DO and High Resolution Spectroscopy of CH_{3}O and HFCO”, Ph.D. Thesis, U.of California, Berkeley, 2001